# FLARE: Feed-forward Geometry, Appearance and Camera Estimation from Uncalibrated Sparse Views

### v3

@@ Abstract
Figure 1: We present flare, a feed-forward system that simultaneously recovers high-quality poses, geometry and appearance from uncalibrated sparse views within 0.5s. Our model excels in the scenarios with a camera circling around the subject, and also shows robust generalization to real-world casual captures, such as an indoor bedroom. In the center below, we casually captured 6 random images of a bedroom with minimal overlap between views. Our method demonstrates high-quality geometry reconstruction even in this challenging case.
@@ 1 Introduction
We present flare, a novel feed-forward and differentiable system that infers high-quality geometry, appearance, and camera parameters from uncalibrated sparse-view images. Direct optimization of these parameters from images often presents significant learning difficulties, frequently converging to sub-optimal solutions with distorted geometry and blurry textures. To address these challenges, we introduce a novel cascade learning paradigm that progressively estimates camera poses, geometry, and appearance, relaxing traditional requirements for 3D reconstruction such as dense image views, accurate camera poses, and wide baselines. Our cascade learning paradigm lies in decomposing the challenging optimization problem into sequential stages, using camera poses as proxies for each stage. The central concept is that a camera pose frames a 2D image within a 3D observation frustum, reducing the learning complexity for subsequent tasks. Our method starts with a neural pose predictor that estimates coarse camera poses from sparse-view images. These initial poses provide geometric cues that facilitate a transformer-based architecture to refine the poses, compute 3D pointmaps, and predict 3D Gaussians for novel-view synthesis. For geometry prediction, we introduce a two-stage approach instead of direct global prediction. First, we estimate camera-centric pointmaps in individual camera coordinates. Then, a neural scene projector unifies these local pointmaps into a coherent global structure. With this approach, we enable faster convergence in geometry learning and reduce geometric distortion for challenging scenes.
@@ 2 Related Work
Pose-free Novel-view Synthesis. Recent research has made significant progress in novel-view synthesis from uncalibrated images. One line of research focuses on jointly optimizing camera poses and radiance fields from dense-view images. BARF [36], NeRF-- [81], and subsequent works [26, 3, 72] have advanced this approach. Several recent methods [15, 30, 14] have extended the 3D representation from NeRF to 3D Gaussians. Another research direction [52, 46, 84, 28] focuses on developing feed-forward novel-view synthesis for unposed images. SRT [53] proposes the first pose- and geometry-free framework for novel view synthesis, while LEAP [27] pioneers pose-free radiance field reconstruction by directly estimating scene geometry and radiance fields. FlowCam [63] and FlowMap [64] introduce 2D flow to enable unsupervised learning of generalizable 3D reconstruction, though their performance degrades in sparse-view settings. PF-LRM [78] estimates camera poses by predicting point maps and solving a differentiable perspective-n-point (PnP) problem, but shows limited generalization to complex 3D scenes. PF3plat [22] achieves coarse alignment of 3D Gaussians by leveraging pre-trained models for monocular depth estimation and visual correspondence. Splatt3R [62] and NopoSplat [94] utilize DUSt3R [80] or MASt3R [32] to predict point maps as proxy geometry and subsequently learn 3D Gaussians for sparse-view reconstruction. However, existing approaches are either restricted to two-view scenarios or produce suboptimal rendering results due to imperfect geometry estimates from DUSt3R or MASt3R. Our work presents a differentiable system that simultaneously optimizes camera parameters, geometry, and appearance, achieving superior generalization across diverse real-world scenes.
@@ 3.1 Neural Pose Predictor
We parametrize each camera pose as a 7-dimensional vector comprising the absolute translation and a normalized quaternion. With this neural pose predictor, we can estimate coarse camera poses as initialization for subsequent geometry and appearance learning. We observe that the estimated poses do not need to be very accurate—only approximating the ground truth distribution is enough. This aligns with our key insight: camera poses, even imperfect, provide essential geometric priors and spatial initialization, which significantly reduces the optimization complexity for geometry and appearance reconstruction.
@@ 3.2 Multi-view Geometry Estimation
Camera-centric Geometry Estimation. Learning geometry in camera space aligns with the image formation process, as each view directly observes local geometry from its perspective. This also simplifies the geometry learning by focusing on local structures visible in each view, rather than reasoning about complex global spatial relationships. We tokenize images into low-resolution tokens and concatenate them with camera tokens derived from coarse pose estimates $\mathcal{P}_{c}$. These tokens are then fed into a transformer architecture $\mathrm{F}_{l}(\cdot)$ to estimate low-resolution 3D point tokens $\mathcal{T}_{l}=\{\mathbf{T}^{local}_{i}\}_{i=1}^{N}$. The self-attention mechanism in the transformer can perform association across different views and exploit the geometric cues from cameras. The 3D point tokens are subsequently fed into a DPT-based [48] decoder $\mathrm{D}_{l}(\cdot)$ for spatial upsampling to obtain dense pointmaps $\mathcal{G}_{l}=\{\mathbf{G}^{\text{local}}_{i}\}_{i=1}^{N}$ and confidence map $\mathcal{C}_{l}=\{\mathbf{C}^{\text{local}}_{i}\}_{i=1}^{N}$ in local camera space. Meanwhile, we further refine the initial camera poses by introducing additional learnable pose tokens $\mathcal{Q}_{f}$ into the network, which output refined pose estimates $\mathcal{P}_{f}$ alongside the geometry prediction:
@@
Global Geometry Projection. We aim to transform camera-centric geometry predictions into a consistent global geometry using refined camera poses. However, this transformation is challenging since imperfect pose estimates make direct geometric reprojection unreliable. Rather than using geometric transformation, we propose a learnable geometry projector $\mathrm{F}_{g}(\cdot)$ that transforms local geometry $\mathcal{G}_{l}$ into global space, conditioned on the estimated poses $\mathcal{C}_{f}$. This learned approach is more robust to pose inaccuracies compared to direct geometric projection. For computational efficiency, we utilize the low-resolution 3D point tokens $\mathcal{T}_{l}$ rather than the dense camera geometry $\mathcal{G}_{l}$ as the input:
@@ 4 Experiment
Implementation details.
Our model is trained from scratch using 8 views as input, without any pre-trained models, except for the encoder. The neural pose predictor $\mathrm{F}_{p}(\cdot)$ consists of 12 transformer blocks with channel width 768. In cascade geometry estimation, our camera-centric geometry estimator $\mathrm{F}_{l}(\cdot)$ and global geometry projector $\mathrm{F}_{g}(\cdot)$ both use 12 transformer blocks with channel width 768, and the DPT-head consists of 4 convolutional blocks to achieve 4$\times$ upsampling. We use the Adam [31] optimizer with an initial learning rate of $1\times 10^{-4}$, gradually decreasing to $1\times 10^{-5}$. We train our model on 64 NVIDIA A800 GPUs for 200 epochs with an input resolution of $512\times 384$. The training takes approximately 14 days to complete. We first train the geometry estimation network and camera pose estimator without the Gaussian regression head for approximately 7 days using a batch size of 8 per GPU. Subsequently, we introduce the Gaussian regression head while keeping the coarse camera pose estimator frozen, training for an additional 7 days with a reduced batch size of 4 per GPU. The loss weights $\lambda_{\text{pose}}$, $\lambda_{\text{geo}}$, and $\lambda_{\text{splat}}$ are set to 5, 1, and 4 by default. More implementation details are presented in the supplementary materials.
Inference.
4.1 Multi-view Pose Estimation
Baseline. For camera pose estimation, we compare our method with recent deep optimization-based methods, including DUSt3R [80], MASt3R [32], VGGSfM [77] and RelPose [98]. We also include traditional SfM methods such as COLMAP [55] and PixSfM [57], as well as feed-forward method PoseDiffusion [76]. Note that among all baselines, PoseDiffusion is trained on the RealEstate10K dataset while others are not.
Comparison. We show quantitative comparison in Tab. 1. Conventional SfM methods like COLMAP and PixSfM rely on time-consuming bundle adjustment, resulting in slow inference speed for camera pose estimation. These optimization-based methods also struggle with sparse views where feature correspondences are difficult to establish. DUSt3R and MASt3R employ global alignment optimization, which not only leads to slow inference but also limits their performance due to their two-view geometry learning paradigm that cannot effectively leverage multi-view associations. Compared to other feed-forward approaches, our method achieves superior performance, thanks to the coarse-to-fine pose estimation strategy and multi-task learning with geometry and appearance.
@@ 4.1 Multi-view Pose Estimation
PoseDiffusion [76] 0.91 0.73 78.2
Ours 0.97 0.74 84.6
4.2 Sparse-view 3D Reconstruction
@@ 5 Discussion
We propose flare, a feed-forward model that can infer high-quality camera poses, geometry, and appearance from sparse-view uncalibrated images within 0.5 seconds. We demonstrate that leveraging camera poses as reference points effectively simplifies complex 3D learning and thus we propose a novel cascade learning paradigm that progressively estimates camera poses, geometry, and appearance, leading to substantial improvements over previous methods. Our model, trained on a set of public datasets, learns strong reconstruction priors that generalize robustly to challenging scenarios, such as very sparse views captured in real-world settings, enabling photo-realistic novel view synthesis.
5.1 Limitations and Future Work.
Our approach cannot reconstruct thin structures perfectly. Additionally, camera pose estimation becomes not very accurate when dealing with out-of-distribution camera trajectories. For future work, we plan to explore advanced multi-scale geometry representations to better handle thin structures, and incorporate out-of-distribution algorithms in pose estimation to improve robustness across diverse camera trajectories.
@@ Acknowledgements
We gratefully acknowledge Tao Xu for his assistance with the evaluation, Yuanbo Xiangli for insightful discussions, and Xingyi He for his Blender visualization code.
@@ Appendix B Experiments
Table 8: Architecture of Neural Pose Predictor.
Layer Name Module Type Details
decoder_embed Linear in_features=1024, out_features=768, bias=True
dec_blocks ModuleList 12 x DecoderBlock
dec_norm LayerNorm shape=(768,), eps=1e-06, elementwise_affine=True
dec_cam_norm LayerNorm shape=(768,), eps=1e-06, elementwise_affine=True
dec_blocks2 ModuleList 12 x DecoderBlock
pose_head CameraPredictor Same as VGGSfM [77]
Table 9: Architecture of Camera-centric Geometry Estimation Network.
Layer Name Module Type Details
decoder_embed Linear in_features=1024, out_features=768, bias=True
pose_token_ref Parameter Initialized with shape (1, 1, 768)
pose_token_source Parameter Initialized with shape (1, 1, 768)
embed_pose MLP Encodes pose embeddings for input poses
cam_cond_encoder ModuleList 12 x Attention Block
cam_cond_embed ModuleList 12 x Linear
adaLN_modulation ModuleList 12 x Sequential (Shift, Scale, Gate Modulation)
dec_blocks ModuleList 12 x DecoderBlock
dec_norm LayerNorm shape=(768,), eps=1e-6, elementwise_affine=True
inject ModuleList 3 x Linear for token injection
enc_inject ModuleList 3 x LayerNorm for token preprocessing
point_map_head DPT_head Same as MASt3R [32]
dec_cam_norm LayerNorm shape=(768,), eps=1e-6, elementwise_affine=True
pose_head CameraPredictor Same as VGGSfM [77]
Table 10: Architecture of Global Geometry Projector.
Layer Name Module Type Details
embed_pose MLP Encodes pose embeddings for input poses
decoder_embed Linear in_features=1024, out_features=768, bias=True
pose_token_ref Parameter Learnable reference token for pose embedding
pose_token_source Parameter Learnable source token for multi-view poses
cam_cond_encoder ModuleList 12 x Attention Block
cam_cond_embed ModuleList 12 x Linear
dec_blocks ModuleList 12 x DecoderBlock
dec_blocks_cross ModuleList 12 x Cross-attention
dec_norm LayerNorm shape=(768,), eps=1e-6, elementwise_affine=True
point_map_head DPT_head Same as MASt3R [32]
Appendix C Network Architecture
We employ the CroCo encoder [82] as our pretrained encoder model, keeping its weights frozen during training. The architecture of our neural pose predictor, trained from scratch, is detailed in Tab. 8. Similarly, our camera-centric geometry estimation network is trained from scratch, and its architecture is also outlined in Tab. 9. Additionally, the architecture of our global geometry projector, trained from scratch, is presented in Tab. 10. The appearance model takes as input the dense features extracted from the VGG and DPT heads. These features are fused using a two-layer U-Net, which combines the information from both sources. The fused features are then processed through separate convolutional layers to predict the final outputs, such as color, opacity, and high-frequency spherical harmonics coefficients.

### v4

@@ Abstract
Figure 1: We present flare, a feed-forward approach that simultaneously recovers high-quality poses, geometry and appearance from uncalibrated sparse views within 0.5s. Our model excels in the scenarios with a camera circling around the subject, and also shows robust generalization to real-world casual captures, such as an indoor bedroom. In the central area below, we casually captured six random bedroom images with minimal overlap. Our method demonstrates high-quality geometry reconstruction even in this challenging case.
@@ 1 Introduction
We present flare, a novel feed-forward and differentiable system that infers high-quality geometry, appearance, and camera parameters from uncalibrated sparse-view images. Direct optimization of these parameters from images often presents significant learning difficulties, frequently converging to sub-optimal solutions with distorted geometry and blurry textures. To address these challenges, we introduce a novel cascade learning paradigm that progressively estimates camera poses, geometry, and appearance, relaxing traditional requirements for 3D reconstruction such as dense image views, accurate camera poses, and wide baselines. Our cascade learning paradigm lies in decomposing the challenging optimization problem into sequential stages, using camera poses as proxies for each stage. The central concept is that a camera pose frames a 2D image within a 3D observation frustum, reducing the learning complexity for subsequent tasks. Our method starts with a neural pose predictor that estimates coarse camera poses from sparse-view images. These initial poses provide geometric cues that facilitate a transformer-based architecture to refine the poses, compute point maps, and predict 3D Gaussians for novel-view synthesis. For geometry prediction, we introduce a two-stage approach instead of direct global prediction. First, we estimate camera-centric point maps in individual camera coordinates. Then, a neural scene projector unifies these local point maps into a coherent global structure. With this approach, we enable faster convergence in geometry learning and reduce geometric distortion for challenging scenes.
@@ 2 Related Work
Pose-free Novel-view Synthesis. Recent research has made significant progress in novel-view synthesis from uncalibrated images. One line of research focuses on jointly optimizing camera poses and radiance fields from dense-view images. BARF [36], NeRF-- [81], and subsequent works [26, 3, 72] have advanced this approach. Several recent methods [15, 30, 14] have extended the 3D representation from NeRF to 3D Gaussians. Another research direction [52, 46, 83, 28] focuses on developing feed-forward novel-view synthesis for unposed images. SRT [53] proposes the first pose- and geometry-free framework for novel view synthesis, while LEAP [27] pioneers pose-free radiance field reconstruction by directly estimating scene geometry and radiance fields. FlowCam [63] and FlowMap [64] introduce 2D flow to enable unsupervised learning of generalizable 3D reconstruction, though their performance degrades in sparse-view settings. PF-LRM [78] estimates camera poses by predicting point maps and solving a differentiable perspective-n-point (PnP) problem, but shows limited generalization to complex 3D scenes. PF3plat [22] achieves coarse alignment of 3D Gaussians by leveraging pre-trained models for monocular depth estimation and visual correspondence. Splatt3R [62] and NopoSplat [93] utilize DUSt3R [80] or MASt3R [32] to predict point maps as proxy geometry and subsequently learn 3D Gaussians for sparse-view reconstruction. However, existing approaches are either restricted to two-view scenarios or produce suboptimal rendering results due to imperfect geometry estimates from DUSt3R or MASt3R. Our work presents a differentiable system that simultaneously predicts camera parameters, geometry, and appearance, achieving superior generalization across diverse real-world scenes.
@@ 3.1 Neural Pose Predictor
We parametrize each camera pose as a 7-dimensional vector comprising the absolute translation and a normalized quaternion. With this neural pose predictor, we can estimate coarse camera poses as initialization for subsequent geometry and appearance learning. We observe that the estimated poses do not need to be very accurate—only approximating the ground truth distribution is enough. This aligns with our key insight: camera poses, even imperfect, provide essential geometric priors and spatial initialization, which significantly reduces the complexity for geometry and appearance reconstruction.
@@ 3.2 Multi-view Geometry Estimation
Camera-centric Geometry Estimation. Learning geometry in camera space aligns with the image formation process, as each view directly observes local geometry from its perspective. This also simplifies the geometry learning by focusing on local structures visible in each view, rather than reasoning about complex global spatial relationships. We tokenize images into image tokens and concatenate them with camera tokens derived from coarse pose estimates $\mathcal{P}_{c}$. These tokens are then fed into a transformer architecture $\mathrm{F}_{l}(\cdot)$ to estimate local point tokens $\mathcal{T}_{l}=\{\mathbf{T}^{local}_{i}\}_{i=1}^{N}$. The self-attention mechanism in the transformer can perform association across different views and exploit the geometric cues from cameras. The local point tokens are subsequently fed into a DPT-based [48] decoder $\mathrm{D}_{l}(\cdot)$ for spatial upsampling to obtain dense point maps $\mathcal{G}_{l}=\{\mathbf{G}^{\text{local}}_{i}\}_{i=1}^{N}$ and confidence map $\mathcal{C}_{l}=\{\mathbf{C}^{\text{local}}_{i}\}_{i=1}^{N}$ in local camera space. Meanwhile, we further refine the initial camera poses by introducing additional learnable pose tokens $\mathcal{Q}_{f}$ into the network, which output refined pose estimates $\mathcal{P}_{f}$ alongside the geometry prediction:
@@
Global Geometry Projection. We aim to transform camera-centric geometry predictions into a consistent global geometry using refined camera poses. However, this transformation is challenging since imperfect pose estimates make direct geometric reprojection unreliable. Rather than using geometric transformation, we propose a learnable geometry projector $\mathrm{F}_{g}(\cdot)$ that transforms local geometry $\mathcal{G}_{l}$ into global space, conditioned on the estimated poses $\mathcal{P}_{f}$. This learned approach is more robust to pose inaccuracies compared to direct geometric projection. For computational efficiency, we utilize the local point tokens $\mathcal{T}_{l}$ rather than the dense camera geometry $\mathcal{G}_{l}$ as the input:
@@ 4 Experiment
PoseDiffusion [76] 0.78 0.27 51.6
Ours 0.97 0.65 79.9
@@
Implementation details.
Our model is trained from scratch using 8 views as input, without any pre-trained models, except for the encoder. The neural pose predictor $\mathrm{F}_{p}(\cdot)$ consists of 12 transformer blocks with channel width 768. In cascade geometry estimation, our camera-centric geometry estimator $\mathrm{F}_{l}(\cdot)$ and global geometry projector $\mathrm{F}_{g}(\cdot)$ both use 12 transformer blocks with channel width 768. We use the Adam [31] optimizer with an initial learning rate of $1\times 10^{-4}$, gradually decreasing to $1\times 10^{-5}$. We train our model on 64 NVIDIA A800 GPUs for 200 epochs with an input resolution of $512\times 384$. The training takes approximately 14 days to complete. More implementation details are presented in the supplementary materials.
Inference.
4.1 Multi-view Pose Estimation
Baseline. For camera pose estimation, we compare our method with recent deep optimization-based methods, including DUSt3R [80], MASt3R [32], VGGSfM [77] and RelPose [97]. We also include traditional SfM methods such as COLMAP [55] and PixSfM [57], as well as feed-forward method PoseDiffusion [76].
Comparison. We show quantitative comparison in Tab. 1. Conventional SfM methods like COLMAP and PixSfM rely on time-consuming bundle adjustment, resulting in slow inference speed for camera pose estimation. These optimization-based methods also struggle with sparse views where feature correspondences are difficult to establish. DUSt3R and MASt3R employ global alignment optimization, which not only leads to slow inference but also limits their performance due to their two-view geometry learning paradigm that cannot effectively leverage multi-view associations. Compared to other feed-forward approaches, our method achieves superior performance, thanks to the coarse-to-fine pose estimation strategy and multi-task learning with geometry.
4.2 Sparse-view 3D Reconstruction
@@ 5 Discussion
We introduce flare, a feed-forward model that can infer high-quality camera poses, geometry, and appearance from sparse-view uncalibrated images within 0.5 seconds. We propose a novel cascade learning paradigm that progressively estimates camera poses, geometry, and appearance, leading to substantial improvements over previous methods. Our model, trained on a set of public datasets, learns strong reconstruction priors that generalize robustly to challenging scenarios, such as very sparse views captured in real-world settings, enabling photo-realistic novel view synthesis.
@@ Acknowledgements
We gratefully acknowledge Tao Xu for his assistance with the evaluation, Yuanbo Xiangli for insightful discussions, and Xingyi He for his Blender visualization code. This work was partially supported by NSFC (No. U24B20154), Zhejiang Provincial Natural Science Foundation of China (No. LR25F020003), Ant Group, and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.